According to a research​ institution, men spent an average of ​$135.62 on​ Valentine's Day gifts in 2009. Assume the standard deviation for this population is ​$40 and that it is normally distributed. A random sample of 10 men who celebrate​ Valentine's Day was selected.

Complete parts a through e.

a. Calculate the standard error of the mean. sigma Subscript x overbarσ=

​(Round to two decimal places as​ needed.)

b. What is the probability that the sample mean will be less than ​$130?

P(x<$130)=

​(Round to four decimal places as​ needed.)

c. What is the probability that the sample mean will be more than ​$140?

P(x>$140)=

​(Round to four decimal places as​ needed.)

d. What is the probability that the sample mean will be between ​$110 and ​$165​?

P($110≤x≤$165)=

​(Round to four decimal places as​ needed.)

e. Identify the symmetrical interval that includes​ 95% of the sample means if the true population mean is ​$135.62.

≤x ≤​

​(Round to the nearest dollar as​ needed.)
Answer #1

Solution :

Given that,

mean = \mu = $135.62

standard deviation = \sigma = ​$40

a ) n = 10

\mu\bar x = 135.62

\sigma\bar x  =  (\sigma /\sqrtn) = ( ​40 / \sqrt10 ) = 12.65

b ) P (  \bar x < 130 )

P ( \bar x - \mu\bar x/\sigma\bar x) < (130 - 135.62 / 12.65)

P ( z < - 5.62 /12.65 )

P ( z < - 0.44 )

Using z table

= 0.3300

Probability = 0.3300

c ) P (  \bar x > 140 )

= 1 - P (  \bar x < 140 )

= 1 - P ( \bar x - \mu\bar x/\sigma\bar x) < (140 - 135.62 / 12.65)

= 1 - P ( z < 4.38 /12.65 )

= 1 - P ( z < 0.35)

Using z table

= 1 - 0.6368

= 0.3632

Probability = 0.3632

d ) ( 110 ≤ \bar x ≤ 165)

P ( 110 - 135.62 / 12.65) ≤ ( \bar x - \mu\bar x/\sigma\bar x)   ≤ ( 165 - 135.62 / 12.65)

P ( - 25.62 / 12.65 ≤ z ≤ 29.38 / 12.65 )

P (- 2.03 ≤ z ≤ 2.32 )

P (z ≤ 2.32 ) - p ( z ≤ - 2.03 )

Using z table

= 0.9898 - 0.0212

= 0.9686

Probability = 0.9686

e ) The critical z-value = 0.05\alpha  = 0.05 is ZC= 1.96 ZC= 1.96.

symmetrical interval = ( \bar x - ( ZC * \sigma/ \sqrtn ) ,  \bar x + ( ZC * \sigma/ \sqrtn )

= ( 135.62 - 1.96 * 40 \sqrt10, 135.62 -+1.96 * 40 \sqrt10)

= (135.62 - 24.792, 135.62 + 24.792

=(110.828,160.412)

= ( 111 , 160 )

$111 ≤ x ≤ $160